# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ 4.3 Coming down from infinity
(i). By Lemma 4.10, we see $\mathcal{M}$ is totally ordered with respect to the stochastic order. Take $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ as the set in (4.13), and Theorem 4.8 (iv) gives $\mathcal{M}\subset K\subset\mathcal{P}_{\infty}(\mathbb{R})$. The finiteness of $\mathcal{M}$ is postponed to the end of the proof.
(ii). By Theorem 4.4, the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow. Theorem 4.8 gives a eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$, and this set $K$ attracts all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Now, apply Theorem 4.2, whose statements (i)(ii) say that, there are two invariant measures $\underline{\nu},\overline{\nu}\in\mathcal{P}_{2}(\mathbb{R})$ such that the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$. By the definition of global attractor, one has $\mathcal{M}\subset\omega(K)$. Obviously, $\underline{\nu}=\inf_{\leq_{\mathrm{st}}}\mathcal{M}$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\mathcal{M}$. Next, we need to show the global convergence to the order interval $[\underline{\nu},\overline{\nu}]$. Since $\omega(K)$ is the global attractor, the inclusion $\omega(K)\subset[\underline{\nu},\overline{\nu}]$ tells us, for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\underline{\nu},\overline{\nu}])\to 0$ as $t\to\infty$.
(iii). This is a consequence of Theorem 4.2 (iii).
@@
Then the number of zeros of $F$ is exactly the number of invariant measures of the equation (1.1), and each zero of $F$ is the mean value of the corresponding invariant measure. We assume there are infinitely many invariant measures, and argue by contradiction. By the statement (ii), the set of invariant measures $\mathcal{M}$ is contained in an order interval $[\underline{\nu},\overline{\nu}]$. So the boundedness of $[\underline{\nu},\overline{\nu}]$ in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 3.4) entails boundedness of the zero set of $F$ in $\mathbb{R}$. Since invariant measures are infinitely many, the zero set of $F$ has an accumulation point. By Assumption 1 (iii)(v), we have $F$ is an analytic function, and this implies $F$ is identically zero. This is impossible, and we complete the proof. ∎
@@
In Remark 2.2, we show that the set of invariant measures cannot attract all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Here, by Theorem 4.2 (i)(ii) and Theorem 4.8 (i), there exists a connected compact invariant set (the global attractor) within $[\underline{\nu},\overline{\nu}]$ which attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
@@
The uniqueness of invariant measure enatils that $\underline{\nu}=\nu=\overline{\nu}$. Then the antisymmetry of stochastic order implies that $[\underline{\nu},\overline{\nu}]=\{\nu\}$. Hence, this corollary is a direct consequence of Theorem 2.1 (ii). ∎
5 Unbounded Open Neighbourhood Attracted by Supremum (Infimum) Invariant Measure
This section attempts to prove Theorem 2.3 and Corollary 2.4. In Section 5.1, we find the number of invariant measures of the equation (1.1) is bounded by the zero-crossing number of the function $V^{\prime}$. In Section 5.2, we show the existence of positively invariant neighbourhoods of invariant measures, for both the semigroup $P_{t}^{*}$ and the measure-iterating map $\Psi$. After providing a structure theorem of local attractors of order-preserving semigroups in Section 5.3, we proceed to prove Theorem 2.3 and Corollary 2.4 by tracing solutions of (1.1) backward along connecting orbits.
5.1 Number of invariant measures of granular media equations
In this subsection, we provide the prior information on the number of invariant measures. Assuming $V^{\prime}$ has $(2n-1)$ simple zeros, Alecio [1] shows that the equation (1.1) has only one invariant measure if $\left\lvert\sigma\right\rvert$ is above a certain upper critical threshold, and has $(2n-1)$ invariant measures if $\left\lvert\sigma\right\rvert$ is below another lower critical threshold. There is a gap between these two critical thresholds, and the information is missing in the intermediate stage. Here, for any non-degenerate $\sigma$, we prove that the number of invariant measures of the equation (1.1) does not exceed the zero-crossing number of $V^{\prime}$ in Propoaition 5.3. This facilitates the quantitative results in our main theorems.
Let $f:\mathbb{R}\to\mathbb{R}$ be a function. Recall that its zero-crossing number $Z(f)$ is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$. It is easy to check that, if $V^{\prime}$ has $(2n-1)$ simple zeros, then $Z(V^{\prime})=2n-1$. Next, we will see the zero-crossing number of the function $V^{\prime}$ is essential to determine the number of invariant measures. We begin with a simple case where the zero-crossing number is one.
@@ 5.1 Number of invariant measures of granular media equations
Now define another function $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
@@
By Lemma 5.1, the conclusion is true for $n=1$, and we prove by induction. Assume the conclusion is true for $n$. If the function $f$ is continuous and $Z(f)=n+1$, then we may assume there is an increasing sequence $-\infty=x_{-1}<x_{0}<\dots<x_{n}<x_{n+1}=\infty$ such that, $f$ changes its sign on any two adjacent intervals $(x_{k-1},x_{k})$, $k=0,1,\dots,n+1$. Define $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
@@
Suppose Assumption 1 holds. If the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has zero-crossing number $Z(V^{\prime})=2n-1$, then the equation (1.1) has at most $(2n-1)$ invariant measures.
@@
Assumption 1 (iv)(v) tell us $\theta>0$ and $\sigma^{2}(x)>0$ for all $x\in\mathbb{R}$, so $Z(V^{\prime}/\theta\sigma^{2})=Z(V^{\prime})=2n-1$. Lemma 5.2 implies that $F$ has at most $(2n-1)$ zeros, and thus, the equation (1.1) has at most $(2n-1)$ invariant measures. ∎
5.2 Positively invariant neighbourhoods of invariant measures
In this subsection, we prove the existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods of invariant measures under $P^{*}$ when assuming the locally dissipative condition. The existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods is also obtained for the measure-iterating map $\Psi$ defined in (5.13).
Theorem 5.4.
Suppose that Assumption 1 holds. If the locally dissipative condition holds at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then the equation (1.1) has an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. Moreover, $B(\delta_{a},r_{a})$ is positively invariant sets, i.e.,
$P_{t}^{*}B(\delta_{a},r_{a})\subset B(\delta_{a},r_{a}),\ \text{ for all }\ t\geq 0.$ (5.3)
Proof.
By (ii) in the definition of local dissipation (Definition 2.3), we know that there exists $0<\tilde{r}_{a}<r_{a}$ such that
$g_{a}(z,r_{a}^{2})>0,\ \text{ for all }z\geq\tilde{r}_{a}^{2}.$ (5.4)
To prove (5.3), it suffices to prove that $\overline{B(\delta_{a},r)}$ is positively invariant for all $\tilde{r}_{a}\leq r<r_{a}$, which implies that $B(\delta_{a},r)$ is positively invariant for all $\tilde{r}_{a}<r\leq r_{a}$.
For any fixed $\tilde{r}_{a}\leq r<r_{a}$ and $\mu\in\overline{B(\delta_{a},r)}$, let
$\mathcal{T}_{\mu}:=\sup\big{\{}t\geq 0:P_{s}^{*}\mu\in\overline{B(\delta_{a},r)}\ \text{ for all }\ 0\leq s\leq t\big{\}}.$
To prove that $\overline{B(\delta_{a},r)}$ is positively invariant, it suffices to show that $\mathcal{T}_{\mu}=\infty$. By Theorem 4.4, we know that $P^{*}:\mathbb{R}\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is jointly continuous, so $\alpha_{t}:=\mathcal{W}_{2}^{2}(P^{*}_{t}\mu,\delta_{a})$ is continuous in $t$. Then it is equivalent to show that
$\mathcal{T}_{\mu}=\sup\big{\{}t\geq 0:\alpha_{s}\leq r^{2}\ \text{ for all }\ 0\leq s\leq t\big{\}}=\infty.$
Otherwise, if $\mathcal{T}_{\mu}<\infty$, then it follows from the continuity of $\alpha$ and definition of $\mathcal{T}_{\mu}$ that there exists $T_{0}\geq\mathcal{T}_{\mu}$ and $\epsilon_{0}>0$ such that
$\alpha_{T_{0}}=r^{2},\ \text{ and }\ r^{2}<\alpha_{t}\leq r_{a}^{2}\ \text{ for all }\ T_{0}<t\leq T_{0}+\epsilon_{0}.$ (5.5)
Let $X_{t}^{\xi}$ be the solution of (1.1) with intial distribution $\mathcal{L}(\xi)=\mu$ and starting time 00. Then $P_{t}^{*}\mu=\mathcal{L}(X_{t}^{\xi})$ for all $t\geq 0$. For any $a\in\mathbb{R}$, denote $\tau^{a}:\mathbb{R}\to\mathbb{R}$ by the translation $x\mapsto x+a$ and $\tau^{a}_{\#}:\mathcal{P}(\mathbb{R})\to\mathcal{P}(\mathbb{R})$ by $\tau^{a}_{\#}\mu=\mu\circ(\tau^{a})^{-1}$. Applying Itô’s formula to $|X_{t}^{\xi}-a|^{2}$ on $[T_{0},T_{0}+\epsilon_{0}]$, we have for any $T_{0}<t\leq T_{0}+\epsilon_{0}$,
$\begin{split}|X_{t}^{\xi}-a|^{2}&=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*P^{*}_{s}\mu)(X_{s}^{\xi})+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*\tau^{-a}_{\#}P^{*}_{s}\mu)(X_{s}^{\xi}-a)+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &\leq|X_{T_{0}}^{\xi}-a|^{2}-\int_{T_{0}}^{t}g_{a}\big{(}|X_{s}^{\xi}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s},\end{split}$ (5.6)
where we have used the fact
$\|\tau^{-a}_{\#}P^{*}_{s}\mu\|_{2}^{2}=\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}P^{*}_{s}\mu(x)=\mathcal{W}_{2}^{2}(P^{*}_{s}\mu,\delta_{a})=\alpha_{s}\in[r^{2},r_{a}^{2}]\ \text{ for all $s\in[T_{0},T_{0}+\varepsilon]$}.$
Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we get
$\alpha_{T_{0}+\epsilon_{0}}\leq r^{2}-\int_{T_{0}}^{T_{0}+\epsilon_{0}}g_{a}(\alpha_{t},r_{a}^{2})\mathop{}\!\mathrm{d}t.$ (5.7)
Then by (5.4), (5.5) and (5.7), we conclude that $\alpha_{T_{0}+\epsilon_{0}}<r^{2}$ which contradicts (5.5). Hence,
$P_{t}^{*}\overline{B(\delta_{a},r)}\subset\overline{B(\delta_{a},r)},\ \text{ for all }\ \tilde{r}\leq r<r_{a},$ (5.8)
and (5.3) follows.
It remains to prove there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. We first give the following two claims.
Claim 1: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$. If $P_{t}^{*}$ is compact for some $t>0$, then there exists a fixed point $\mu_{t}$ of $P_{t}^{*}$ in $S$, i.e., $P^{*}_{t}\mu_{t}=\mu_{t}$.
Claim 2: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$, and $P_{t}^{*}$ is compact for all $t>0$, then there exists an invariant measure $\mu$ in $S$, i.e., $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
Proof of Claim 1: Let $S_{t}=P_{t}^{*}S$ and $\overline{\mathrm{Conv}(S_{t})}$ be the closed convex hull of $S_{t}$. Since $P_{t}^{*}$ is compact, it follows $S_{t}$ is relatively compact. Noting also that $S$ is closed convex positively invariant, we have $\overline{\mathrm{Conv}(S_{t})}\subset S$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and
$P_{t}^{*}\overline{\mathrm{Conv}(S_{t})}\subset P_{t}^{*}S=S_{t}\subset\overline{\mathrm{Conv}(S_{t})}.$
Note that $\overline{\mathrm{Conv}(S_{t})}$ being compact convex in $\mathcal{P}_{2}(\mathbb{R})$ entails that it is compact convex in $\mathcal{P}_{1}(\mathbb{R})$. Let us consider the space $\mathcal{M}_{1}(\mathbb{R})$ of finite signed Borel measures on $\mathbb{R}$ with finite first moments,
$\mathcal{M}_{1}(\mathbb{R}):=\left\{\mu\in\mathcal{M}(\mathbb{R}):\int_{\mathbb{R}}\left\lvert x\right\rvert\mathop{}\!\mathrm{d}\left\lvert\mu\right\rvert(x)<\infty\right\}$
with the following norm:
[…]
where a $1$-Lipschitz function $f$ means $|f(x)-f(y)|\leq|x-y|$ for all $x,y\in\mathbb{R}$. By [48, Lemma 3.4 and Remark 3.2], we have the following properties of $\mathcal{M}_{1}(\mathbb{R})$:
$(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed vector space; (5.10)
$\mathcal{P}_{1}(\mathbb{R})$ is a closed convex subset of $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$. (5.11)
Hence, $\overline{\mathrm{Conv}(S_{t})}$ is a compact convex subset of a normed vector space. Then the Tychonoff fixed-point theorem yields that there exists $\mu_{t}\in\overline{\mathrm{Conv}(S_{t})}\subset S$ such that $P^{*}_{t}\mu_{t}=\mu_{t}$.
Proof of Claim 2: Fix a positive sequence $\{t_{n}\}_{n\geq 1}$ such that $t_{n}\downarrow 0$ as $n\to\infty$. For any $n\geq 1$, consider the continuous compact map $P^{*}_{t_{n}}:S\to S$. Then Claim 1 yields that there exists $\mu_{n}\in S$ such that $P^{*}_{t_{n}}\mu_{n}=\mu_{n}$. Let $k_{n}$ be an integer such that $k_{n}t_{n}>t_{1}$ for all $n\geq 1$. Then
$\mu_{n}=P_{k_{n}t_{n}}^{*}\mu_{n}=P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}\mu_{n})\in P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}S)\subset P_{t_{1}}^{*}S\subset S\ \text{ for all $n\geq 1$}.$
Since $P_{t_{1}}^{*}S\subset S$ is relatively compact and $S$ is closed, there exists a subsequence of $\{\mu_{n}\}_{n\geq 1}$, which still denotes by $\{\mu_{n}\}_{n\geq 1}$, and a $\mu\in S$ such that $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$. Then we show that $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
In fact, for any $t>0$, we have $t=k_{n,t}t_{n}+\tau_{n,t}$, $k_{n,t}\in\mathbb{N}$ and $\tau_{n,t}\in[0,t_{n})$. Then,
$\mathcal{W}_{2}(P^{*}_{t}\mu,\mu)$ $\leq\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{t}\mu_{n},\mu)$
$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}(P^{*}_{k_{n,t}t_{n}}\mu_{n}),\mu)$
$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu).$
Since $\tau_{n,t}\to 0$ and $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$, by virtue of the joint continuity of the $P^{*}$, we have $\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})\to 0$ and $\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu)\to 0$ as $n\to\infty$. Hence, we obtain that $P^{*}_{t}\mu=\mu$ and the claim is proved.
Now we are in a position to show that there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. Since $P^{*}_{t}$ is compact for all $t>0$ (see Theorem 4.8) and $\overline{B(\delta_{a},\tilde{r}_{a})}$ is bounded closed convex positively invariant under $P^{*}$, by Claim 2, there exists an invariant measure $\mu\in\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a})$. Moreover, Lemma 4.7 shows that $\mu\in\mathcal{P}_{\infty}(\mathbb{R})$. Now, we conclude that the invariant measure $\mu\in B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. ∎
Under Assumption 1, we know that SDE (5.12) has a unique invariant measure $\rho^{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ (see e.g., Meyn-Tweedie [54, 55] or Lemma 4.9 (ii)). Define a measure-iterating map $\Psi:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by
[…]
Then we have the following theorem.
Theorem 5.5.
Proof.
Note that there is some $\tilde{r}_{a}\in(0,r_{a})$ such that (5.4) holds. Now fix any $\mu\in\overline{B(\delta_{a},r_{a})}$. Let $\{X_{t}^{\mu,x}\}_{t\geq 0}$ be the solution of (5.12) with initial value $x\in\mathbb{R}$ and starting time 0. Similar to (5.6), applying Itô’s formula to $|X_{t}^{\mu,a}-a|^{2}$ yields
$|X_{t}^{\mu,a}-a|^{2}\leq-\int_{0}^{t}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{0}^{t}(X_{s}^{\mu,a}-a)\sigma(X_{s}^{\mu,a})\mathop{}\!\mathrm{d}B_{s}.$
Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we have that for any $t>0$,
$g_{a}\bigg{(}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s,r_{a}^{2}\bigg{)}\leq\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\big{]}\mathop{}\!\mathrm{d}s\leq-\frac{1}{t}\mathbf{E}[|X_{t}^{\mu,a}-a|^{2}]\leq 0.$ (5.14)
According to Lemma 4.9 (ii), we know that there exist $C_{\mu}>0,\lambda_{\mu}>0$ such that for any $t\geq 0$,
$d_{2}\big{(}\mathcal{L}(X_{t}^{\mu,a}),\Psi(\mu)\big{)}\leq C_{\mu}(1+|a|^{2})e^{-\lambda_{\mu}t},$
where
$d_{2}(\nu_{1},\nu_{2})=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\nu_{1}-\nu_{2}|(x),$ (5.15)
Then
$\big{|}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}-\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\big{|}=\Big{|}\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}\big{(}\mathcal{L}(X_{s}^{\mu,a})-\Psi(\mu)(x)\big{)}\Big{|}\leq 2C_{\mu}(1+|a|^{2})^{2}e^{-\lambda_{\mu}s}.$
Thus
$\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s=\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}.$
Then the continuity of $g_{a}(\cdot,r_{a}^{2})$ and (5.14) yield $g_{a}(\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2},r_{a}^{2})\leq 0$. Hence, by (5.4), $\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\leq\tilde{r}_{a}^{2}$, i.e., $\Psi(\mu)\in\overline{B(\delta_{a},\tilde{r}_{a})}$. Hence,
$\Psi\big{(}\overline{B(\delta_{a},r_{a})}\big{)}\subset\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a}).$
The proof is completed. ∎
Remark 5.1.
Note that $\mu$ is an invariant measure of the equation (1.1) if and only if $\Psi(\mu)=\mu$, and then Theorem 5.5 indicates that there is no invariant measure on $\partial\overline{B(\delta_{a},r_{a})}:=\overline{B(\delta_{a},r_{a})}\setminus B(\delta_{a},r_{a})$.

### v2

@@ 3.2 Coming down from infinity
(i). By Lemma 3.8, we see $\mathcal{M}$ is totally ordered with respect to the stochastic order. Take $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ as the set in (3.8), and Proposition 3.6 (iv) gives $\mathcal{M}\subset K\subset\mathcal{P}_{\infty}(\mathbb{R})$. The finiteness of $\mathcal{M}$ is postponed to the end of the proof.
(ii). By [36, Theorem 3.5], the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow. Proposition 3.6 gives a eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$, and this set $K$ attracts all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Now, apply Proposition 3.2, whose statements (i)(ii) say that, there are two invariant measures $\underline{\nu},\overline{\nu}\in\mathcal{P}_{2}(\mathbb{R})$ such that the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$. By the definition of global attractor, one has $\mathcal{M}\subset\omega(K)$. Obviously, $\underline{\nu}=\inf_{\leq_{\mathrm{st}}}\mathcal{M}$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\mathcal{M}$. Next, we need to show the global convergence to the order interval $\underline{\nu},\overline{\nu}]$. Since $\omega(K)$ is the global attractor, the inclusion $\omega(K)\subset[\underline{\nu},\overline{\nu}]$ tells us, for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\underline{\nu},\overline{\nu}])\to 0$ as $t\to\infty$.
(iii). This is a consequence of Proposition 3.2 (iii).
@@
Then the number of zeros of $F$ is exactly the number of invariant measures of the equation (1.1), and each zero of $F$ is the mean value of the corresponding invariant measure. We assume there are infinitely many invariant measures, and argue by contradiction. By the statement (ii), the set of invariant measures $\mathcal{M}$ is contained in an order interval $\underline{\nu},\overline{\nu}]$. So the boundedness of $\underline{\nu},\overline{\nu}]$ in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 2.4) entails boundedness of the zero set of $F$ in $\mathbb{R}$. Since invariant measures are infinitely many, the zero set of $F$ has an accumulation point. By Assumption 1 (iii)(v), we have $F$ is an analytic function, and this implies $F$ is identically zero. This is impossible, and we complete the proof. ∎
@@
In Remark 1.3, we show that the set of invariant measures cannot attract all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Here, by Proposition 3.2 (i)(ii) and Proposition 3.6 (i), there exists a connected compact invariant set (the global attractor) within $\underline{\nu},\overline{\nu}]$ which attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
@@
The uniqueness of invariant measure enatils that $\underline{\nu}=\nu=\overline{\nu}$. Then the antisymmetry of stochastic order implies that $\underline{\nu},\overline{\nu}]=\{\nu\}$. Hence, this corollary is a direct consequence of Theorem 1.2 (ii). ∎
4 Basins of attraction
This section attempts to prove Theorem 1.1, Theorem 1.4, Corollary 1.5, Theorem 1.6, Theorem 1.7 and Theorem 1.8. In Section 4.1, we prove that the number of invariant measures of the equation (1.1) is bounded by the zero-crossing number of the function $V^{\prime}$. Using the fact that solutions can be traced backward along connecting orbits, we establish a local attractor theorem in Section 4.2 to show that the forward limit of a connecting orbit is locally attracting. Then, we finish the proof of our main results, especially the verification of the local dissipative condition under the assumptions of Theorem 1.1.
4.1 Number of invariant measures
In this subsection, we provide the prior information on the number of invariant measures. We show the number of invariant measures is at most the zero-crossing number of the derivative $V^{\prime}$ of confining potential, regardless of the ranges of $\theta$ and $\sigma$ (see Proposition 4.3). This helps to make our main theorems quantitative (Theorem 1.6-1.8), and overcomes difficulties caused by qualitative results and missing parameter in the previous literature on phase transitions (see e.g., [1, 52]). One the other hand, this proposition allows $V^{\prime}$ vanishes on some interval (Theorem 1.8), which is out of the settings of many existing works. Also, it provides an insight that the essential factor in the maximal number of invariant measures is the zero-crossing number of $V^{\prime}$, instead of the number of simple zeros of $V^{\prime}$.
Let $f:\mathbb{R}\to\mathbb{R}$ be a function. Recall that its zero-crossing number $Z(f)$ is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$. It is easy to check that, if $V^{\prime}$ has (2n1)(2n-1) simple zeros, then $Z(V^{\prime})=2n-1$. Next, we will see the zero-crossing number of the function $V^{\prime}$ is essential to determine the number of invariant measures. We begin with a simple case where the zero-crossing number is one.
@@ 4.1 Number of invariant measures
Now define another function $\tilde{F}$$\mathbb{R}\to\mathbb{R}$ by
@@
By Lemma 4.1, the conclusion is true for $n=1$, and we prove by induction. Assume the conclusion is true for $n$. If the function $f$ is continuous and $Z(f)=n+1$, then we may assume there is an increasing sequence $\infty=x_{-1}<x_{0}<\dots<x_{n}<x_{n+1}=\infty$ such that, $f$ changes its sign on any two adjacent intervals $(x_{k-1},x_{k})$, $k=0,1,\dots,n+1$. Define $\tilde{F}$$\mathbb{R}\to\mathbb{R}$ by
@@
Suppose Assumption 1 holds. If the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has zero-crossing number $Z(V^{\prime})=2n-1$, then the equation (1.1) has at most (2n1)(2n-1) invariant measures.
@@
Assumption 1 (iv)(v) tell us $\theta>0$ and $\sigma^{2}(x)>0$ for all $x\in\mathbb{R}$, so $Z(V^{\prime}/\theta\sigma^{2})=Z(V^{\prime})=2n-1$. Lemma 4.2 implies that $F$ has at most (2n1)(2n-1) zeros, and thus, the equation (1.1) has at most (2n1)(2n-1) invariant measures. ∎